CONDITION NUMBER BOUNDS FOR PROBLEMS WITH INTEGER COEFFICIENTS

GREGORIO MALAJOVICH

Abstract. An apriori bound for the condition number associated to each of the following problems is given: general linear equation solving, minimum squares, nonsymmetric eigenvalue problems, solving univariate polynomials, solving systems of multivariate polynomials. It is assumed that the input has integer coefficients and is not on the degenerate locus of the respective problem (i.e. the condition number is finite). Then condition numbers are bounded in terms of the dimension and of the bit-size of the input.

In the same setting, bounds are given for the speed of convergence of the following iterative algorithms: QR without shift for the symmetric eigenvalue problem, and Graeffe iteration for univariate polynomials.